OpenStudy (anonymous):

Use a system of equations to solve the following problem: Charles High School has a total of 94 boys and girls who play sports. If the number of boys is 22 more than twice the number of girls, how many boys play sports at this high school? 70 48 24 80

OpenStudy (kropot72):

Let number of girls = g Let number of boys = b b + g = 94 g = 94 - b ..............(1) b = 2g + 22 ...........(2) Substituting the value for g in equation (1) into equation (2) gives b = 2(94 - b) + 22 .........(3) Expanding (3) gives b = 188 - 2b + 22 .........(4) Can you now solve (4) to find the value of b?